In how many ways can 6 people be seated if two people insisted of sitting next to each other?
Answer by Edwin McCravy(20054)

Let the people be A,B,C,D,E,F 

Suppose A and B insist on sitting next to each other

Then we have 

these 5 things to arrange in a line: 

AB,C,D,E,F 

which is 5! ways, 

plus 

these 5 things to arrange in a line:

BA,C,D,E,F 

which is also 5! ways.

Answer 2×5! = 2×120 = 240 ways. 

Edwin